# The role of porosity in the transition to inertial regime in porous media flows

### v1

@@ 3 Results
where the first term on the right-hand side represents the vortex volume normalized by the total volume of fluid and the other one – the total velocity magnitude per unit recirculation zone volume. We consider each of those two components separately. Fig. 8 shows the visualization of the percolating volumes and recirculation zones for low- and high-porosity SC samples, respectively, for two different Reynolds numbers each. In both systems, as $Re_{k}$ rises, the recirculation zones grow in volume, which in the case of low-porosity system is visible through the percolating volume shrinking to a straight jet connecting the inlet and the outlet. To quantify the changes in the recirculation zones size, we plot their volume, $I[1]_{v}$, normalized by the total volume of fluid, $I[1]$, in Fig. 10. In the low-porosity system, there already exist recirculation zones in the Darcy regime. The increase in their volume starts at $Re_{k}\approx 10^{-3}$, which corresponds to the start of $T_{\Omega}$ decrease in Fig. 9. The vortex volume then continues to grow until the end of the steady state regime, with significant slowdown for $Re_{k}>4\cdot 10^{-2}$, the value at which the slope of the $T_{\Omega}$ rise in Fig. 9 decreases. This reduction in the recirculation volume growth pace indicates the strong geometrical confinement for the vortex development. The total increase in the volume of the recirculation zone is about 46% of the fluid volume, or – equivalently – $46/90\approx 0.51$ of the volume of the solid obstacles. On the other hand, in the high-porosity system there is no recirculation in the Darcy regime. The increase of its volume starts at $Re_{k}\approx 2\cdot 10^{2}$, which is about two decades later than the start of $T_{\Omega}$ decrease for this system in Fig. 9. As the recirculation volume starts to grow, it does so with a slightly increasing slope until the end of the steady-state regime. The increase in the volume of the recirculation zone in this case is about $0.63$ of the solid obstacle volume. The growth of the recirculation zones in the case of high-porosity SC systems was highlighted by other authors [17] as the reason for lower drag compared to the flow over a single obstacle.
@@
Origin of the $T_{\Omega,v}$ decrease and rise – energy confinement in the recirculation oznes
@@
The increase of the recirculation zone volume can explain the decrease of $T_{\Omega}$ because it directly implies the rectification of all the percolating streamlines and thus the decrease of $T_{s}$ in Eq. 14. However, it fails to address the rise of the volume-based tortuosity at the highest Reynolds numbers. This is why now we turn our attention to the integral of the velocity magnitude inside the vortex per vortex volume, $I[V]_{v}/I[1]_{v}$ (further in the text we shall refer to this quantity as the specific velocity magnitude in vortex, for brevity). Its dependence on the Reynolds number is shown in Fig. 11 for both SC systems. We note that since in Darcy regime there was no recirculation zone present in the higher-porosity SC system, we assumed the value of the specific velocity magnitude in vortex to be zero in those cases. For both SC systems, the specific velocity magnitude in vortex increases with Reynolds number. In the low-porosity sample, the rise proceeds from the lowest $Re_{k}$ with the first order slope ($p=1$) and experiences a jump in the slope around $Re_{k}=10^{-2}$. It later attains the slope of the order approximately one again, in the range of $Re_{k}$ corresponding to those where the growth of $T_{\Omega}$ slows down (cf. Fig. 9). The initial first-order dependence is explained by the linearity of the Stokes equation and proportionality between the body force magnitude $g$ and the mean flow velocity $\langle\boldsymbol{V}\rangle$ in the Darcy regime (i.e., when $g$ is increased, the velocity field inside the vortex is only scaled proportionally by a constant). The subsequently higher slope indicates that a non-trivial mechanism involving the momentum transfer to the recirculation zone and the growth of its volume comes into play in the inertial regime. At the same time, an analogy between the Darcy regime and the inertial regime above $Re_{k}=3\cdot 10^{-2}$ can be drawn in a sense that both are characterized by the recirculation zone growing more in kinetic energy, rather than volume. This further highlights the importance of the geometrical confinement in the emergence of inertial effects in low-porosity samples [14]. In the high-porosity system, the start of the rapid growth of the specific velocity magnitude in vortex happens roughly at $Re_{k}$ corresponding to the start of the growth of the recirculation zone volume (cf. Fig. 10). It is interesting that even though the integral of the velocity magnitude over the whole fluid volume, $I[V]$, achieved at the highest Reynolds number is about $36$ times larger in the high-porosity system, the specific velocity magnitude in vortex is over 2 times higher in low-porosity system ($\approx 140$ vs. $\approx 60$). This suggests that the strong geometrical confinement leads to the entrapment of relatively large amounts of the kinetic energy in the recirculation zones, quantified also by other authors using the variance of the kinetic energy [11, 13, 12, 14].
@@
In the end, in the low-porosity stochastic and SC systems, the greater kinetic energy confinement in the recirculation zones and the cease of their volume growth at high Reynolds numbers lead to the more pronounced increase of $T_{\Omega}$, compared to the higher-porosity counterparts.
@@ 4 Conclusions
The presented findings on the differences in behavior of the streamline-based and volume-integrated tortuosity during the Darcy-inertial transition advance the understanding of the mechanisms behind the inertia onset in the studied porous systems. We show that porosity and the associated geometrical confinement are two of the key factors determining the behavior of the system in the inertial regime. The growth of the recirculation zones volume and the corresponding rectification of the percolating streamlines are captured by both types of tortuosity via their decrease in the initial stage of inertia onset. The later dynamics of the inertial effects emergence is strongly dependent on the system’s geometry. In low-porosity media, the geometrical confinement prevents the recirculation zones from growing freely in volume, thus mainly the kinetic energy confined within them increases. This is represented by the saturation of the streamline-based tortuosity at its minimal value, while the volume-integrated tortuosity grows rapidly in this regime. This is because the volume-integrated tortuosity contains information about the whole fluid volume, while the streamline-based one – only about the percolating volume. On the other hand, in the high-porosity systems the growth of the recirculation zones volume is not limited by the solid walls. Due to this, the increase of the volume-integrated tortuosity is less pronounced compared to the low-porosity samples, and the range of Reynolds numbers over which the two definitions of tortuosity are compliant with one another can extend almost to the end of the steady-state flow regime.
@@
The investigated physical phenomena are closely related to the tools used to quantify them. The presented results show the utility of the two discussed measures of the inertial effects in porous media – streamline-based and volume-integrated tortuosity. Each of them separately carries information about various forms of the inertial effects in the flow – $T_{s}$ is more sensitive to the changes occurring solely in the percolating volume, while in $T_{\Omega}$ the recirculation zones contribution is also visible. The difference between the values of the two, as well as the magnitude of the growth of $T_{\Omega}$, can be treated as an indicator of the confinement of the recirculation zones. The presented results also highlight which aspects of the inertia onset in porous media are better captured by which definition of tortuosity and why, which will allow for a more sensitive choice between the two in more application-oriented problems. One may stipulate, for instance, that the information about the recirculation zones carried by the volume-integrated tortuosity might prove useful when the mixing and reaction phenomena in the porous sample are of interest. On the other hand, the exclusive contribution of the percolating volume to the streamline-based tortuosity may carry more valuable information for studies concerned with the mass transport through the sample, e.g. during gas extraction, where the gaseous phase is often transported through preferential channels and the residuals of the fluid phase occupy the dead-end pores and cavities.
@@ Appendix A Relation to other measures for the onset of inertia
To give a better perspective on our present findings, for the discussed stochastic and simple cubic systems we calculate other the measures of the inertia onset. Fig. 12 shows the relation between the mean flow velocity in the $x_{0}$-direction versus the magnitude of the forcing term. A comparison with Figs. 4 and  9 shows that for the higher-porosity samples of each type the Reynolds number at which the mean velocity–forcing term relation becomes nonlinear corresponds roughly to the beginning of the fall of the tortuosities ($Re_{k}\approx 0.52$ and $5.75$ for the stochastic and simple cubic system, respectively). For the lower-porosity samples, the transition to the nonlinear dependence occurs later, around the minimum of $T_{\Omega}$ ($Re_{k}\approx 0.56$ and $9.38\cdot 10^{-3}$, for the stochastic and simple cubic system, respectively). Note also that for the low-porosity simple cubic sample, the divergence from linearity in $\langle V_{0}\rangle$$g$ relation is visible the least from among all four samples. This is in contrast to the $T_{\Omega}$, which for the same sample experiences the biggest jump in value between the Darcy regime and the highest $Re_{k}$. This suggests that the volume-integrated tortuosity might be a good indicator for the onset of inertia for extremely low-porosity systems.
Figure 12: Mean flow velocity in the $x_{0}$-direction versus the magnitude of the forcing term for the stochastic systems (left) and simple cubic systems (right). The dashed lines present the first-order scaling slopes anchored at the data points with the lowest value of $g$ (for some systems those points are not visible at the presented axes ranges). The values at the arrows show the $Re_{k}$ for the data points around which the relation becomes non-linear. The acceleration and velocity for stochastic systems are given in lattice units.
@@
$f_{c,0}=-\left(\frac{\langle\boldsymbol{V}\rangle}{|\langle\boldsymbol{V}\rangle|}-\frac{k_{0}\boldsymbol{g}}{\mu|\langle\boldsymbol{V}\rangle|}\right)\cdot\hat{\boldsymbol{x}}_{0}=-1+\frac{k_{0}g}{\mu\langle V_{0}\rangle}$ (16)
@@
on the Reynolds number $Re_{k}$ for the studied stochastic and simple cubic systems. In Eq. (16), $\hat{\boldsymbol{x}}_{0}$ is the unit vector aligned with the $x_{0}$-axis and the last equality holds under the assumption of the mean flow aligned with the $x_{0}$ direction, namely $\langle\boldsymbol{V}\rangle=\langle V_{0}\rangle\hat{\boldsymbol{x}}_{0}$. The weak inertia regime is characterized by the $f_{c,0}\propto Re_{k}^{2}$ scaling, while the strong inertia regime – by $f_{c,0}\propto Re_{k}$ scaling. In the case of stochastic samples, both regimes are vaguely visible, which suggest that the transition between the two regimes takes place over the majority of the considered Reynolds numbers. In the case of simple cubic systems, the weak inertia regime is clearly visible for both porosities. For the lower-porosity sample, in the lowest $Re_{k}$ cases the scaling becomes below-second order, which was also observed by other authors [15]. The deviation from the weak inertia regime apperas approximately at $Re_{k}=10^{-2}$, for which the $T_{\Omega}$ reaches its minimum (cf. Fig 9). For the highest Reynolds numbers, the relation changes into sublinear, however it endures over a short range of $Re_{k}$ before the steady-state regime ends. It is possible that one observes the regime above strong inertia where $f_{c,0}\propto Re_{k}^{2}+Re_{k}$. One might argue though that this regime is related to the reduction in the increase of the slope of the volume-based tortuosity at the highest $Re_{k}$. For the higher-porosity simple cubic sample, relation becomes sub-quadratic around $Re_{k}=1$, the value at which the tortuosities in Fig. 9 begin to decrease. Also in this case, for the highest Reynolds numbers, the $f_{c,0}(Re_{k})$ relation is clearly non-linear. However, the highest-Reynolds number results should be regarded with caution as one operates close to the limit of the stability of the steady-state solver.
@@
Figure 13: The inertial corection $f_{c,0}$ (Eq. (16)) versus the Reynolds number for the stochastic systems (left) and simple cubic systems (right). The dashed lines present the first- and second-order scaling slopes for the stochastic samples and $1/2$– and second order scaling slopes for the simple cubic samples. The inset in the right subplot shows the magnification of the highest-$Re_{k}$ region for the lower-porosity sample.

### v2

@@ 3.4 Origin of the $T_{\Omega}$ decrease and rise – volume of the recirculation zones
where the first term on the right-hand side represents the vortex volume normalized by the total volume of fluid, and the other one – the total velocity magnitude per unit recirculation zone volume. We consider each of those two components separately. Figs. 17 and 18 show the visualization of the percolating volumes and recirculation zones for low- and high-porosity SC samples, respectively, for two different Reynolds numbers each. In both systems, as $Re_{k}$ rises, the recirculation zones grow in volume, which in the case of the low-porosity system is visible through the percolating volume shrinking to a straight jet connecting the inlet and the outlet. To quantify the changes in the size of the recirculation zones, we plot their volume, $I[1]_{v}$, normalized by the total volume of fluid, $I[1]$, in Fig. 19. In the low-porosity system, there already exist recirculation zones in the Darcy regime (in the channels perpendicular to the main flow, cf. left visualization in Fig. 14). The increase in their volume starts at $Re_{k}\approx 5\cdot 10^{-4}$, which corresponds to the start of $T_{\Omega}$ decrease in Fig. 16. The vortex volume then continues to grow until the end of the steady state regime, with significant slowdown for $Re_{k}>10^{-2}$, the value at which the slope of the $T_{\Omega}$ rise in Fig. 16 decreases. This reduction in the recirculation volume growth pace indicates the strong geometrical confinement for the vortex development. We highlight that at the highest $Re_{k}$, there already exist secondary vortices in the perpendicular channels (cf. right visualization in Fig. 14). The total increase in the volume of the recirculation zone is about 56% of the fluid volume, or – equivalently – $0.56\cdot 0.1/0.9\approx 0.062$ of the volume of the solid obstacles. On the other hand, in the high-porosity system, there is no recirculation in the Darcy regime (cf. left visualization in Fig. 15). The increase of its volume starts at $Re_{k}\approx 2\cdot 10^{2}$, which is about two decades later than the visible start of $T_{\Omega}$ decrease for this system in Fig. 16. This discrepancy can be related to the ‘formation of inertial cores’ [44] when the inertia already starts to change the velocity field, although no separation yet occurs. As the recirculation volume starts to grow, it does so with an increasing slope until the end of the steady-state regime (cf. right visualization in Fig. 15). The increase in the volume of the recirculation zone in this case is about $0.63$ of the solid obstacle volume. The growth of the recirculation zones in the case of high-porosity SC systems was highlighted by other authors [19] as the reason for lower drag compared to the flow over a single obstacle. Finally, we note that as the considered SC systems have a regular and simple geometry, phenomena like the non-strictly-monotonic growth of the volume of the recirculation zones [45] related to the instantaneous reorganization of the percolating flow paths are not visible here. Nevertheless, the start of the recirculation zones volume growth can still be related to the deviation from Darcy’s law at $Re\approx 10^{-3}$ (compare the data for $\phi=0.1$ SC system in Figs. 19 and 21). Interestingly, no high-wavenumber oscillations can be seen in the values of $T_{\Omega}(Re_{k})$ in the case of the stochastic samples (cf. Fig. 11), which suggests that the reorganization of the percolating flow paths might not impact the velocity integrals of the volume-integrated tortuosity.
Figure 17: Visualizations of the percolating volumes in $\phi=0.1$ SC system for the same values of $Re_{k}$ as in Fig. 14. The gray surfaces are the walls of the obstacles, and only a few chosen walls are shown for clarity. The visualizations use the histograms from which the percolating volumes were found. We show only the percolating volumes for clarity.
Figure 18: Visualizations of the recirculating volumes for $\phi=0.999$ SC system for the same values of $Re_{k}$ as in Fig. 15. The gray surfaces are the walls of the obstacles. The visualizations use the histograms from which the recirculation volumes were found. We show only the recirculation volumes for clarity.
@@
3.5 Origin of the $T_{\Omega}$ decrease and rise – energy confinement in the recirculation zones
@@
The increase of the recirculation zone volume can explain the decrease of $T_{\Omega}$ because it directly implies the straightening of all the percolating streamlines and thus the decrease of $T_{s}$ in Eq. 14. However, it fails to address the rise of the volume-based tortuosity at the highest Reynolds numbers. This is why now we turn our attention to the integral of the velocity magnitude inside the vortex per vortex volume, $I[V]_{v}/I[1]_{v}$ (further in the text, we shall refer to this quantity as the specific velocity magnitude in vortex, for brevity). Its dependence on the Reynolds number is shown in Fig. 20 for both SC systems. We note that since there is no recirculation zone present in the higher-porosity SC system in the Darcy regime, we assume that the value of the specific velocity magnitude in the vortex is zero in those cases. For both SC systems, the specific velocity magnitude in vortices increases with Reynolds number. In the low-porosity sample, the rise proceeds from the lowest $Re_{k}$ with the first-order slope ($p=1$) and experiences a jump in the slope around $Re_{k}=3\cdot 10^{-3}$. It later attains the slope of the order approximately one again, in the range of $Re_{k}$ corresponding to those where the growth of $T_{\Omega}$ slows down (cf. Fig. 16). The initial first-order dependence is explained by the linearity of the Stokes equation and proportionality between the body force magnitude $g$ and the mean flow velocity $\langle\boldsymbol{V}\rangle$ in the Darcy regime (i.e., when $g$ is increased, the velocity field inside the vortex is only scaled proportionally by a constant). The subsequently higher slope indicates that a non-trivial mechanism involving the momentum transfer to the recirculation zone and the growth of its volume comes into play in the inertial regime. At the same time, an analogy between the Darcy regime and the inertial regime above $Re_{k}=10^{-2}$ can be drawn in the sense that both are characterized by the recirculation zone growing more in kinetic energy than volume. This further highlights the importance of the geometrical confinement in the emergence of inertial effects in low-porosity samples [16]. In the high-porosity system, the start of the rapid growth of the specific velocity magnitude in the vortex happens roughly at $Re_{k}$ when the recirculation zone volume starts to grow (cf. Fig. 19). It is interesting that even though the integral of the velocity magnitude over the whole fluid volume, $I[V]$, achieved at the highest Reynolds number is about $40$ times larger in the high-porosity system (results not shown in the plots), the specific velocity magnitude in vortex is about 2 times higher in low-porosity system ($\approx 140$ vs. $\approx 70$). This suggests that strong geometrical confinement leads to the accumulation of relatively large amounts of the kinetic energy in the recirculation zones, quantified also by other authors using the variance of the kinetic energy [13, 15, 14, 16].
@@
In the end, in the low-porosity stochastic and SC systems, the greater kinetic energy confinement in the recirculation zones and the cessation of their volume growth at high Reynolds numbers lead to the more pronounced increase of $T_{\Omega}$, compared to the higher-porosity counterparts.
@@ 4 Conclusions
The presented findings on the differences in behavior of the streamline-based and volume-integrated tortuosity during the Darcy-inertial transition in simple cubic porous media advance the understanding of the mechanisms behind the inertia onset in the studied porous systems. We show that porosity and the associated geometrical confinement are two of the key factors determining the behavior of the system in the inertial regime. The growth of the volume of the recirculation zones and the corresponding straightening of the percolating streamlines are captured by both types of tortuosity via their decrease in the initial stage of inertia onset. The later dynamics of the emergence of the inertial effects are strongly dependent on the system’s geometry. In low-porosity media, the geometrical confinement prevents the recirculation zones from growing freely in volume, thus mainly the kinetic energy confined within them increases. This is represented by the saturation of the streamline-based tortuosity at its minimal value, while the volume-integrated tortuosity grows rapidly in this regime. This is because the volume-integrated tortuosity contains information about the whole fluid volume, while the streamline-based one – only about the percolating volume. On the other hand, in high-porosity systems, the growth of the volume of the recirculation zones is not limited by the solid walls. Due to this, the increase of the volume-integrated tortuosity is less pronounced compared to the low-porosity samples, and the range of Reynolds numbers over which the two definitions of tortuosity are compliant with one another can extend almost to the end of the steady-state flow regime.
@@
The investigated physical phenomena are closely related to the tools used to quantify them. The presented results show the utility of the two discussed measures of the inertial effects in porous media – streamline-based and volume-integrated tortuosity. Each of them separately carries information about various forms of the inertial effects in the flow – $T_{s}$ is more sensitive to the changes occurring solely in the percolating volume, while in $T_{\Omega}$ the contribution of the recirculation zones is also visible. The difference between the values of the two, as well as the magnitude of the growth of $T_{\Omega}$, can be treated as an indicator of the confinement of the recirculation zones. The presented results also highlight which aspects of the inertia onset in porous media are better captured by which definition of tortuosity and why, allowing for a more sensitive choice between the two in more application-oriented problems. One may stipulate, for instance, that the information about the recirculation zones carried by the volume-integrated tortuosity might prove useful when the mixing and reaction phenomena in the porous sample are of interest. On the other hand, the exclusive contribution of the percolating volume to the streamline-based tortuosity may carry more valuable information for studies concerned with the mass transport through the sample, e.g., during gas extraction, where the gaseous phase is often transported through preferential channels and the residuals of the fluid phase occupy the dead-end pores and cavities.
@@ Appendix A Relation to other measures for the onset of inertia
To give a better perspective on our present findings, for the discussed stochastic and simple cubic systems, we additionally calculate other measures of the inertia onset. Fig. 21 shows the relation between the mean flow velocity in the $x_{0}$-direction versus the magnitude of the forcing term (the Darcy’s law, Eq. (1)). A comparison with Figs. 11 and  16 shows that for the higher-porosity samples of each type the Reynolds number at which the mean velocity–forcing term relation becomes nonlinear corresponds roughly to the beginning of the fall of the tortuosities ($Re_{k}\approx 0.49$ and $5.75$ for the stochastic and simple cubic system, respectively). For the lower-porosity samples, the transition to the nonlinear dependence occurs later, around the minimum of $T_{\Omega}$ ($Re_{k}\approx 0.48$ and $1.45\cdot 10^{-3}$, for the stochastic and simple cubic system, respectively). Note also that for the simple cubic samples, the divergence from linearity in $\langle V_{0}\ranglegg$ relation is weaker than in the stochastic systems. This is especially in contrast to the behavior of $T_{\Omega}$ in the low-porosity SC sample, which experiences the biggest jump in value between the Darcy regime and the highest $Re_{k}$ from among all the studied systems. This suggests that the volume-integrated tortuosity might be a good indicator for the onset of inertia for extremely low-porosity systems.
Figure 21: Mean flow velocity in the $x_{0}$-direction versus the magnitude of the forcing term for the stochastic systems (left) and simple cubic systems (right). The dashed lines present the first-order scaling slopes anchored at the data points with the lowest value of $g$ (for some systems, those points are not visible at the presented axes ranges). The values at the arrows show the $Re_{k}$ for the data points around which the relation becomes non-linear. The acceleration and velocity for stochastic systems are given in lattice units.
@@
$f_{c,0}=-\left(\frac{\langle\boldsymbol{V}\rangle}{|\langle\boldsymbol{V}\rangle|\phi}-\frac{k_{0}\boldsymbol{g}}{\mu|\langle\boldsymbol{V}\rangle|}\right)\cdot\hat{\boldsymbol{x}}_{0}=-1+\frac{k_{0}g}{\mu\langle V_{0}\rangle\phi}$ (16)
@@
on the Reynolds number $Re_{k}$ for the studied stochastic and simple cubic systems. In Eq. (16), $\hat{\boldsymbol{x}}_{0}$ is the unit vector aligned with the $x_{0}$-axis, and the last equality holds under the assumption of the mean flow aligned with the $x_{0}$ direction, namely $\langle\boldsymbol{V}\rangle=\langle V_{0}\rangle\hat{\boldsymbol{x}}_{0}$. The weak inertia regime is characterized by the $f_{c,0}\propto Re_{k}^{2}$ scaling, while the strong inertia regime – by $f_{c,0}\propto Re_{k}$ scaling. In the case of stochastic samples, the latter regime is vaguely visible, which suggests that the transition between the two regimes takes place over the majority of the considered Reynolds numbers. In the case of simple cubic systems, the weak inertia regime is clearly visible for both porosities. In the several lowest $Re_{k}$ cases, the scaling becomes below-second order, which was also observed by other authors [17]. In the lower-porosity sample, the deviation from the weak inertia regime appears approximately at $Re_{k}=4\cdot 10^{-3}$, for which the $T_{\Omega}$ reaches its minimum (cf. Fig 16). For the highest Reynolds numbers, the relation changes into sublinear; however, it endures over a short range of $Re_{k}$ before the steady-state regime ends. It is possible that one observes the regime above strong inertia where $f_{c,0}\propto Re_{k}^{2}+Re_{k}$. One may relate this regime to the reduction in the increase of the slope of the volume-based tortuosity at $Re_{k}>2\cdot 10^{-2}$. For the higher-porosity simple cubic sample, the relation becomes sub-quadratic around $Re_{k}=5$, the value at which the tortuosities in Fig. 16 begin to rapidly decrease. Also in this case, for the highest Reynolds numbers, the $f_{c,0}(Re_{k})$ relation is clearly non-linear.
@@
Figure 22: The inertial corection $f_{c,0}$ (Eq. (16)) versus the Reynolds number for the stochastic systems (left) and simple cubic systems (right). The dashed lines present the first- and second-order scaling slopes for the stochastic samples and 1/21/2– and second-order scaling slopes for the simple cubic samples. The inset in the right subplot shows the magnification of the highest-$Re_{k}$ region for the lower-porosity sample.
Appendix B Discretization independence study
To assess the discretization independence of $T_{\Omega}$ for the SC systems, we calculate its values from the solution of the governing equations obtained at different space discretization parameters ($h$ for uniform meshes and $h_{\text{min}}$ for refined meshes, cf. Section 2.1.3). For each porosity, the highest $Re_{k}$ is considered. In the case of the higher-porosity sample, we use the refined discretization and solve the transient problem. The results are presented in Table 1. The change of $T_{\Omega}$ between the subsequent discretizations is calculated relative to the value at the finer discretization.
$\phi=0.1$, $g=3.5\cdot 10^{6}$
$h$ $T_{\Omega}-1$ $\%$ change
$1/115$ $0.279$
$1/150$ $0.255$ $9.444$
$\mathbf{1/195}$ $\mathbf{0.296}$ $\mathbf{13.697}$
$1/253$ $0.299$ $0.996$
$1/330$ $0.301$ $0.822$
$\phi=0.999$, $g=5\cdot 10^{3}$
$h_{\text{min}}$ $\Delta t$ $T_{\Omega}-1$ $\%$ change
$1/116$ $2.25\cdot 10^{-6}$ $8.193\cdot 10^{-5}$
$\mathbf{1/150}$ $\mathbf{2.25\cdot 10^{-6}}$ $\mathbf{8.778\cdot 10^{-5}}$ $\mathbf{6.665}$
$1/196$ $2.25\cdot 10^{-6}$ $9.035\cdot 10^{-5}$ $2.851$
$1/254$ $1.5\cdot 10^{-6}$ $9.137\cdot 10^{-5}$ $1.112$
Table 1: Values above unity of the volume-integrated tortuosity $T_{\Omega}-1$ obtained for the highest forcing term $g$ for SC systems at various space discretization parameters used for the solution of the governing equations. The relative change in each row (rightmost column) is normalized by the value of $T_{\Omega}-1$ from the row above (one coarser discretization). The bold case row indicates the values discussed in the work.
We additionally perform the study of the independence of the $T_{s}$ value on the spacing $h_{s}$ and timestep value $\Delta t$ of the streamline generation procedure in SC samples. In this case, we take a look at the samples of the lowest $Re_{k}$, as these are the ones where the streamlines are expected to be the most tortuous, thus introducing larger errors at the segments of high curvature. We use the velocity fields obtained on the uniform $h=1/150$ or refined $h_{\text{min}}=1/150$ discretizations described in Sec. 2.1.3. We scale streamline seeding distance $h_{s}$ and the timestep for the integration along characteristics $\Delta t_{s}$ by a factor of $2$ between subsequent setups. The results are presented in Table 2
$\phi=0.1$, $Re_{k}=2.61\cdot 10^{-7}$
$h_{s}$ $\Delta t_{s}$ $T_{s}-1$ $\%$ change
 $4\cdot 10^{-3}$ $4\cdot 10^{-3}$ $0.0255$
$2\cdot 10^{-3}$ $2\cdot 10^{-3}$ $0.0244$ $4.598$
$\mathbf{10^{-3}}$ $\mathbf{10^{-3}}$ $\mathbf{0.0240}$ $\mathbf{1.676}$
$5\cdot 10^{-4}$ $5\cdot 10^{-4}$ $0.0238$ $0.861$
$\phi=0.999$, $Re_{k}=5.93\cdot 10^{-11}$
$h_{s}$ $\Delta t_{s}$ $T_{s}-1$ $\%$ change
$2\cdot 10^{-2}$ $2\cdot 10^{-2}$ $7.190\cdot 10^{-4}$
$\mathbf{10^{-2}}$ $\mathbf{10^{-2}}$ $\mathbf{7.343\cdot 10^{-4}}$ $\mathbf{2.085}$
$5\cdot 10^{-3}$ $5\cdot 10^{-3}$ $7.385\cdot 10^{-4}$ $0.561$
$2.5\cdot 10^{-3}$ $2.5\cdot 10^{-3}$ $7.404\cdot 10^{-4}$ $0.254$
Table 2: Values above unity of the volume-integrated tortuosity $T_{s}-1$ obtained for the lowest $Re_{k}$ for SC systems at various discretization parameters used for the streamline generation. The relative change in each row (rightmost column) is normalized by the value of $T_{\Omega}-1$ from the row above (one coarser discretization). The bold case row indicates the values discussed in the work.